Celeste wants to have her hair cut and perm and also go to lunch.She knows she will need $50. The perm costs twice as much as her haircut and she needs $5 for lunch.How much does the perm cost?
Answer by Miceandwine(15)

Okay, so let's set up this problem:
2x(the perm)+x(the haircut)+5(lunch)=50
2x+x+5=50
So the first thing we can do is subtract the 5 from both sides and add the x's
3x+5-5=50-5
3x=45 Now we can divide the whole problem by 3 to get:
x=15
Great, now we know how much the haircut cost! Just double that to find out how much the perm costs, $30. Let's check our answer:
$15(haircut)+$30 (perm)+$5(lunch)=$50 (total)
I hope I helped!